# Nearly forgotten results in development of physical cosmology

### v2

@@ 6 Relikt-1 and COBE discoveries
In 1983, in the Soviet Union the Relikt-1 experiment was conducted aboard the Prognoz-9 spacecraft in order to investigate CMB radiation from space for the first time in history. As many other Prognoz missions, the scientific payload was prepared by the Space Research Institute of the Soviet Academy of Sciences. Igor Strukov was the principal investigator of the Relikt-1 project. The spacecraft Prognoz-9, had an 8 mm band radiometer with an extremely high sensitivity of 35 mK per second and it was launched into a high apogee orbit with a 400,000 km semi-axis. The high orbit was a great advantage of the mission since it allows to reduce of geomagnetic field impact on measurements. A disadvantage of the experiment was that the observations were conducted only in one spectral band, therefore, it was a freedom for a theoretical interpretation of the results, in contrast, multi-band measurements provide very small room for alternative explanations of anisotropy. The radiometer scanned the entire celestial sphere for six months. Computer facilities (and therefore, data analyzing) were relatively slow in the time. Preliminary analysis of anisotropy studies indicate upper limits on anisotropy [72], but in this case even the negative results (upper limits) were extremely important to evaluate a sensitivity to design detectors for next missions, including COBE (COsmic Background Explorer).
At the NASA cite https://lambda.gsfc.nasa.gov/product/relikt/ it is given a brief information about Soviet Relikt-1 mission: ”Prognoz 9, launched on 1 July 1983 into a high-apogee (700,000 km) orbit, included the Relikt-1 experiment to investigate the anisotropy of the CMB at 37 GHz, using a Dicke-type modulation radiometer. During 1983 and 1984 some 15 million individual measurements were made (with 10% near the galactic plane providing some 5000 measurements per point). The entire sky was observed in 6 months. The angular resolution was 5.5 degrees, with a temperature resolution of 0.6 mK. The galactic microwave flux was measured and the CMB dipole observed. A quadrupole moment was found between 17 and 95 microkelvin rms, with 90% confidence level. A map of most of the sky at 37 GHz is available.” After that at this cite references at key Relikt-1 publications are given where information about the Relikt-1 discovery was presented [73, 74]
@@
The discovery of anisotropy by the Relikt-1 spacecraft was first reported officially in January 1992 at the Moscow astrophysical seminar. Relikt-1 team submitted their paper in Soviet Astronomy Letters and Monthly Notices of Royal Astronomical Society on January 19, 1992 and on February 3, 1992, respectively. On April 21, 1992, G. Smoot (the head of DMR experiment aboard the COBE mission) and his co-authors submitted a paper in Astrophysical Journal (Letters) [75]. COBE orbit was rather low (around 1000 km) and therefore, it was a hard problem to separate a signal and a noise which was very strong at the low orbit. On April 22, 1992, Smoot reported at a press conference about the discovery of the CMB anisotropy with the COBE satellite. After that mass media reported this results as the main science success. In 1992, COBE results about discovery of the CMB anisotropy were reported elsewhere. However, no doubt, the COBE collaboration knew results of Relikt-1 team and even quoted their upper limit on the quadrupole anisotropy published at the paper [72]. However, summarizing, one could say that since papers of the Relikt-1 team were submitted on January 19, 1992 and February 3, 1992 in Soviet Astronomy and Monthly Notices of Royal Astronomy Society respectively, but COBE paper was submitted on April 21, 1992, one would conclude that the discovery one quadrupole anisotropy was done by the Relikt-1 team and published in papers [73, 74] and also independently and almost simultaneously the phenomenon was discovered by the COBE-DMR team. In 2006 J. Mather and G. Smoot got Nobel prizes in physics for SMB studies since the COBE-FIRAS group measured CMB spectrum while COBE-DMR group found quadrupole anisotropy. The results of the Relikt-1 group were criticized by the members of the COBE group, but subsequent analysis showed that the results of the Relikt-1 group were in agreement with the results of the WMAP group [76].
@@
Until the end of the 1990s, theorists and astronomers thought that Einstein’s $\Lambda$-term is very small and they adopted $\Lambda=0$. In 1998 astronomers discovered accelerated expansion of the Universe and in this case it is necessary to include into consideration $\Lambda$-term or so called dark energy. However, we have to remember that in 1965 Erast Borisovich Gliner considered cosmological (inflationary) models with accelerating expansion [77, 78] (see also an essay on the scientific life of this remarkable scientist and personality in [79]). Such Gliner’s models looked very exotic and some outstanding Soviet cosmologists (including Ya. B. Zeldovich) criticized them; on the other hand, V. L. Ginzburg and A. D. Sakharov tried to support Gliner and his studies [79, 80]. For instance, recalling his work on cosmology in the 1960s, A. D. Sakharov wrote that ”Gliner wrote about the same thing independently and with greater certainty in the same years” [81]. However, in spite of the undoubted importance of his studies, Gliner’s results were nearly forgotten for decades [79].
@@ 8 Conclusion
This fall marks the 100th anniversary of the death of A. A. Friedmann and the 120th anniversary of the GR creation. The above are some results in the development of gravity and cosmology that are not covered in great detail in the Russian literature, for example, when discussing the Big Bang model in Russian literature, Lemaître’s contribution is often forgotten, and the discovery of Shmaonov and the results of the Relikt-1 mission related to the discovery of the quadrupole anisotropy of the universe are very rarely mentioned. Unfortunately, very often our compatriots are not active to promote achievements of other researchers (including our domestic achievements). In these circumstances people remember the M. I. Monastyrsky’s dictum [82] ”…By the way, we (in Russia) are very fond of lamenting the underestimation of Russian scientists in the West. Russian (Soviet) scientists are the ones who hinder the recognition of Russian (Soviet) scientists the most, however, a careful analysis of the real facts shows…” The discussions of the fact that the quadrupole anisotropy of CMB was detected simultaneously from the analysis of Relikt-1 and COBE data were presented in [83, 84]. Some additional information related to the subject of this article can be found in historical introductions in the recent works [85, 86].
@@ Acknowledgements
The author thanks the organizers of the XXVIth International Seminar on High Energy Physics Problems ”Relativistic Nuclear Physics and Quantum Chromodynamics” (JINR, Dubna) for the invitation to present my contribution as an invited talk at this event and their attention to the research presented in the paper. The author appreciates A. A. Baldin, A. D. Chernin, A. D. Kaminker, V. O. Soloviev, M. V. Tokarev, D. Ya. Yakovlev for useful discussions of the subject.

### v3

@@ 6 Relikt-1 and COBE discoveries
In 1983, in the Soviet Union the Relikt-1161616In Soviet Union the CMB radiation was called the relic (in Russian it sounds like Relikt) radiation. experiment was conducted aboard the Prognoz-9 spacecraft in order to investigate CMB radiation from space for the first time in history. As many other Prognoz missions, the scientific payload was prepared by the Space Research Institute of the Soviet Academy of Sciences. Igor Strukov was the principal investigator of the Relikt-1 project. The spacecraft Prognoz-9, had an 8 mm band radiometer with an extremely high sensitivity of 35 mK per second and it was launched into a high apogee orbit with a 400,000 km semi-axis. The high orbit was a great advantage of the mission since it allows to reduce of geomagnetic field impact on measurements. A disadvantage of the experiment was that the observations were conducted only in one spectral band, therefore, it was a freedom for a theoretical interpretation of the results, in contrast, multi-band measurements provide very small room for alternative explanations of anisotropy. The radiometer scanned the entire celestial sphere for six months. Computer facilities (and therefore, data analyzing) were relatively weak in the time in Russia. Preliminary analysis of anisotropy studies indicate upper limits on anisotropy [75], but in this case even the negative results (upper limits) were extremely important to evaluate a sensitivity to design detectors for next missions, including COBE (COsmic Background Explorer).
At the NASA cite https://lambda.gsfc.nasa.gov/product/relikt/ it is given a brief information about Soviet Relikt-1 mission: ’’Prognoz 9, launched on 1 July 1983 into a high-apogee (700,000 km) orbit, included the Relikt-1 experiment to investigate the anisotropy of the CMB at 37 GHz, using a Dicke-type modulation radiometer. During 1983 and 1984 some 15 million individual measurements were made (with 10% near the galactic plane providing some 5000 measurements per point). The entire sky was observed in 6 months. The angular resolution was 5.5 degrees, with a temperature resolution of 0.6 mK. The galactic microwave flux was measured and the CMB dipole observed. A quadrupole moment was found between 17 and 95 microkelvin rms, with 90% confidence level. A map of most of the sky at 37 GHz is available.’’ After that at this cite references at key Relikt-1 publications are given where information about the Relikt-1 discovery was presented [76, 77].
@@
The discovery of anisotropy by the Relikt-1 spacecraft was first reported officially in January 1992 at the Moscow astrophysical seminar. Relikt-1 team submitted their paper in Soviet Astronomy Letters and Monthly Notices of Royal Astronomical Society on January 19, 1992 and on February 3, 1992, respectively. On April 21, 1992, G. Smoot (the head of DMR experiment aboard the COBE mission) and his co-authors submitted a paper in Astrophysical Journal (Letters) [78]. COBE orbit was rather low (around 1000 km) and therefore, it was a hard problem to separate a signal and a noise which was very strong at the low orbit. On April 22, 1992, Smoot reported at a press conference about the discovery of the CMB anisotropy with the COBE satellite. After that mass media reported these results as the main science success. In 1992, COBE results about discovery of the CMB anisotropy were reported elsewhere. However, no doubt, the COBE collaboration knew results of Relikt-1 team and even quoted their upper limit on the quadrupole anisotropy published at the paper [75]. However, summarizing, one could say that since papers of the Relikt-1 team were submitted on January 19, 1992 and February 3, 1992 in Soviet Astronomy and Monthly Notices of Royal Astronomy Society respectively. COBE paper was submitted on April 21, 1992, one would conclude that the discovery one quadrupole anisotropy was done by the Relikt-1 team and published in papers [76, 77] and also independently and almost simultaneously the phenomenon was discovered by the COBE-DMR team. In 2006 J. Mather and G. Smoot got Nobel prizes in physics for СMB studies since the COBE-FIRAS group measured CMB spectrum while COBE-DMR group found the quadrupole anisotropy. The results of the Relikt-1 group were criticized by the members of the COBE group, but subsequent analysis showed that the results of the Relikt-1 group were in agreement with the results of the WMAP group [79].
@@
Until the end of the 1990s, theorists and astronomers thought that Einstein’s $\Lambda$-term is very small and they adopted $\Lambda=0$. In 1998 astronomers discovered accelerated expansion of the Universe and in this case it is necessary to include into consideration $\Lambda$-term or so called dark energy. However, we have to remember that in 1965 Erast Borisovich Gliner considered cosmological (inflationary) models with accelerating expansion [80, 81] (see also an essay on the scientific life of this remarkable scientist and personality in [82]). Such Gliner’s models looked very exotic and some outstanding Soviet cosmologists (including Ya. B. Zeldovich) criticized them; on the other hand, V. L. Ginzburg and A. D. Sakharov tried to support Gliner and to promote his studies [82, 83]. For instance, recalling his work on cosmology in the 1960s, A. D. Sakharov wrote that ’’Gliner wrote about the same thing independently and with greater certainty in the same years’’ [84]. However, in spite of the undoubted importance of his studies, Gliner’s results were nearly forgotten for decades [82].
@@ 8 Conclusion
This fall marks the 100th anniversary of the death of A. A. Friedmann and the 110th anniversary of the GR creation. The above are some results in the development of gravity and cosmology that are not covered in great detail in the Russian literature, for example, when discussing the Big Bang model in Russian literature, Lemaître’s contribution is often forgotten, and the discovery of Shmaonov and the results of the Relikt-1 mission related to the discovery of the quadrupole anisotropy of the universe are very rarely mentioned. Unfortunately, very often our compatriots are not active to promote achievements of other researchers (including our domestic achievements). In these circumstances people remember the M. I. Monastyrsky’s dictum [85] ’’…By the way, we (in Russia) are very fond of lamenting the underestimation of Russian scientists in the West. Russian (Soviet) scientists are the ones who hinder the recognition of Russian (Soviet) scientists the most, however, a careful analysis of the real facts shows…’’ The discussions of the fact that the quadrupole anisotropy of CMB was detected simultaneously from the analysis of Relikt-1 and COBE data were presented in [86, 87]. Some additional information related to the subject of this article can be found in historical introductions in the recent works [88, 89].
@@ Acknowledgements
The author thanks the organizers of the XXVIth International Seminar on High Energy Physics Problems "Relativistic Nuclear Physics and Quantum Chromodynamics" (JINR, Dubna) for the invitation to present my contribution as an invited talk at this event and their attention to the research presented in the paper. The author appreciates A. A. Baldin, A. D. Chernin, E. E. Donets, A. D. Kaminker, V. O. Soloviev, M. V. Tokarev, D. Ya. Yakovlev for useful discussions of the subject. The author is grateful to the anonymous referee for constructive comments that helped to improve the text of the article.
Funding
This work was supported by ongoing institutional funding. No additional grants to carry out or direct this particular research were obtained.
Conflict of interest
The author declares that he has no conflicts of interest.